# SMILE: SeMantic Ids Enhanced CoLd Item Representation for Click-through Rate Prediction in E-commerce SEarch

### v1

@@ Abstract.
With the rise of modern search and recommendation platforms, insufficient collaborative information of cold-start items exacerbates the Matthew effect of existing platform items, challenging platform diversity and becoming a longstanding issue. Existing methods align items’ side content with collaborative information to transfer collaborative signals from high-popularity items to cold-start items. However, these methods fail to account for the asymmetry between collaboration and content, nor the fine-grained differences among items. To address these issues, we propose SMILE, an item representation enhancement approach based on fused alignment of semantic IDs. Specifically, we use RQ-OPQ encoding to quantize item content and collaborative information, followed by a two-step alignment: RQ encoding transfers shared collaborative signals across items, while OPQ encoding learns items’ differentiated information. Comprehensive offline experiments on large-scale industrial datasets demonstrate SMILE’s superiority, and rigorous online A/B tests confirm statistically significant improvements: item CTR +1.66%, buyers +1.57%, and order volume +2.17%.
@@ 1. Introduction
Existing cold-start solutions fall into two paradigms. Generator-based methods (Chen et al., 2022; Luo et al., 2025; Kong et al., 2024; Vartak et al., 2017) synthesize cold-item representations via warm-item signals: meta-learning (Luo et al., 2025) pre-trains transferable generators for few-sample adaptation, GANs (Chen et al., 2022) align cold embeddings with warm signal distributions, and VAEs (Kong et al., 2024) sample from learned warm-data distributions. Knowledge alignment-based methods (Zhuang et al., 2025; Xu et al., 2024, 2025; Wei et al., 2021) bridge content and collaborative signals: contrastive learning optimizes (Wei et al., 2021) content encoders to match warm collaborative representations, while knowledge distillation (Zhuang et al., 2025) uses content features as a medium to transfer teacher-model (warm-item) knowledge to student models (cold-item).
@@
To alleviate these issues, this paper proposes a novel item representation fusion and enhancement method based on semantic IDs, termed SMILE. Specifically, inspired by OneSearch (Chen et al., 2025): a recently deployed end-to-end generative framework that has achieved successful application in e-commerce search. We leverage its proposed RQ-OPQ encoding (endowed with rich hierarchical semantic and collaborative signals) to enhance cold-start item representations. First, for the coarser-grained RQ encoding, we design an adaptive transfer mechanism to enable bidirectional transfer between high-frequency collaborative signals embedded in item IDs and dense semantic information within RQ encodings. Second, for the residual vectors of items obtained via RQ hierarchical quantization (i.e., OPQ encodings), which encapsulate the discriminative information of different items, we integrate such discriminative information into item representations through contrastive learning. Ultimately, the resulting item representations not only incorporate abundant content and collaborative information but also accurately capture the uniqueness of individual items, as illustrate in Figure 1 (c). The main contributions are as follows:
@@
  • We conduct comprehensive offline experiments and rigorous online A/B test, whose results fully validate the superiority of our proposed SMILE method.
@@ 2.1. Cold Start for Search
Typically, the core of the cold-start problem lies in exploring how to predict user interaction behaviors for new items. In this work, we focus on the click-through rate (CTR) prediction task and investigate how to improve the performance of cold-start tasks through the synergy between Semantic IDs generated by large search models and collaborative signals. In this section, we first formally define the cold-start problem, then elaborate on the proposed SMILE framework in detail. Formally, given a user U (user profile $U_{p}$), the user’s current search query Q, and the user’s historical behavior sequence $H_{U}$, cross features $C$, each sample in the CTR prediction task can be represented as $I=(id,X)$, which includes a label $y\in\{0,1\}$ (indicating whether the sample is clicked by the user): id denotes the random hash ID of the item, and $X$ represents the multi-dimensional feature set of the item (including content features, metadata features, etc.). The goal of cold-start CTR estimation is to learn a discriminative model $f$ to output the click probability of the sample:
@@ 2.3. Semantic Ids Enhanced Item Representation
In this section, we elaborate in detail on SMILE, a cold-start item representation enhancement framework based on semantic IDs. This framework comprises two core modules: the first is an adaptive transfer and alignment mechanism between RQ encoding and item IDs, which aims to learn coarse-grained representations with both semantic information and collaborative signals; the second is an item discriminative information learning module based on OPQ encoding, through which fine-grained discriminative information is injected into item representations, as illustrated in Figure 2.
Figure 2. The framework of SMILE.
@@ 3.2.3. Baselines
Since our proposed SMILE framework is based on semantic ids to enhance the representation of cold start items, we have selected some models that are also based on semantic ids or cold start as our baseline:
SPM_SID (Singh et al., 2024): hashing sub fragments of semantic ids sequences to adaptively obtain the semantics of different items.
DAS (Ye et al., 2025): Through joint training of semantic collaborative models, use contrastive learning to synchronously optimize the quantization and alignment process of semantic ids and item id.
SaviorRec (Yao et al., 2025): A cold start model based on multimodal information and semantic ids collaborative training to enhance item representation.
@@ 3.3. Offline Performance
Table 1. The offline performance of SMILE.
@@
SMILE 0.8725 0.6394 0.8741 0.6405 0.8528 0.6301
To verify the effectiveness of SMILE in the click-through rate (CTR) prediction task, this section compares the performance of SMILE with existing state-of-the-art methods (SPM_SID, DAS, SaviorRec) through offline experiments, as shown in Table 1. In the full-sample (All) scenario, SMILE achieves the best performance in both AUC (+0.38pp) and GAUC (+0.34pp). This proves that SMILE’s overall ranking ability and resistance to user bias are significantly superior to baselines. In the warm item scenario, SMILE still maintains a leading advantage, which indicates that even in scenarios with sufficient user interaction data, SMILE can still optimize the representation quality of warm items and further improve prediction accuracy through the fusion of collaborative signals and semantic information in RQ-OPQ encodings. The cold-start item scenario is the core advantage area of SMILE, The improvement margin is much higher than that in the full-sample and warm item scenarios. This result verifies the effectiveness of SMILE’s design—by leveraging the discriminative information learning of OPQ encoding and the adaptive transfer of RQ encoding, SMILE effectively solves the problem of scarce collaborative signals for cold-start items, making it the current State-of-the-Art (SOTA) method for cold-start CTR prediction tasks.
@@ 3.4. Ablation Study
Table 2. The ablation study resluts of SMILE.
@@
SMILE 0.8725 0.6394 0.8741 0.6405 0.8528 0.6301
to verify the function of each module in SMILE, we conducted sufficient ablation experiments, and the experimental results are shown in the Table 2. Firstly, we directly replace the original random ID in base model by RQ-OPQ encodings (short for only sid). Although it achieves comparable overall performance to the base model, it exhibits inferior performance on warm items compared to the original model due to collaborative signal misalignment. Secondly, we attempted to preserve random IDs and treat SIDs as a type of ID feature (short for iid+SID). This variant has a slight improvement compared to the base model, as the introduction of SID brings additional information increment, but it is still far inferior to the SMILE model. Finally, we ablated the two modules of SMILE (short for iid+RQ, iid+OPQ). Both variants can achieve significant benefits compared to the base model. By integrating the advantages of these two variants, SMILE simultaneously learns the discriminative information and convergent semantic information of items, ultimately achieving the optimal performance.
@@ 3.5. Online A/B Testing
Table 3. The Online A/B testing resluts of SMILE.
@@
SMILE +1.720% +2.230% +3.512% +9.639%
To verify the efficiency of SMILE’s online platform, we compared it with an online ranking model in the e-commerce mall search scenario through rigorous A/B testing. We conducted a 14 day experiment under real online traffic, with 5 days being the AA period and the remaining 9 days being the AB period. The base group and exp group were randomly assigned 7% of the overall real online traffic. The experimental results are shown in the Table 3. Our experimental focus indicators are the buyers and order volume, and we also report the relevant indicators in the cold start scenario in the Table 3. Experiments have shown that SMILE can increase the overall number of buyers by 1.72% and the order volume by 2.23%. Simultaneously, SMILE significantly improves the cold start scenario (buyer number +3.512% and order volume +9.639%) in mall ecommerce search.
@@ 4. Conclusion
To address cold-start challenges in e-commerce, we propose SMILE: it fuses item ID collaboration and semantics RQ via adaptive transfer, and enhances item differentiation with OPQ contrastive learning. Offline experiments and online A/B testing confirm its effectiveness (improving buyer +1.72%, order volume +2.23%). In future work, we will explore the fusion of multimodal semantics and collaborative signals to adapt to more complex cold-start scenarios.

### v3

# COINS: SeMantiC Ids Enhanced COld Item RepresentatioN for Click-through Rate Prediction in E-commerce Search

@@ Abstract.
With the rise of modern search and recommendation platforms, insufficient collaborative information of cold-start items exacerbates the Matthew effect of existing platform items, challenging platform diversity and becoming a longstanding issue. Existing methods align items’ side content with collaborative information to transfer collaborative signals from high-popularity items to cold-start items. However, these methods fail to account for the asymmetry between collaboration and content, nor the fine-grained differences among items. To address these issues, we propose COINS, an item representation enhancement approach based on fused alignment of semantic IDs. Specifically, we use RQ-OPQ encoding to quantize item content and collaborative information, followed by a two-step alignment: RQ encoding transfers shared collaborative signals across items, while OPQ encoding learns items’ differentiated information. Comprehensive offline experiments on large-scale industrial datasets demonstrate COINS’s superiority, and rigorous online A/B tests confirm statistically significant improvements: item CTR +1.66%, buyers +1.57%, and order volume +2.17%.
@@ 1. Introduction
Existing cold-start solutions fall into two paradigms. Generator-based methods (10.1145/3477495.3531897; 10.1145/3690624.3709336; 10.1007/978-981-97-2262-4_9; 10.5555/3295222.3295434) synthesize cold-item representations via warm-item signals: meta-learning (10.1145/3690624.3709336) pre-trains transferable generators for few-sample adaptation, GANs (10.1145/3477495.3531897) align cold embeddings with warm signal distributions, and VAEs (10.1007/978-981-97-2262-4_9) sample from learned warm-data distributions. Knowledge alignment-based methods (10.1145/3696410.3714852; 10.1145/3626772.3657839; 10.1145/3711858; 10.1145/3474085.3475665) bridge content and collaborative signals: contrastive learning optimizes (10.1145/3474085.3475665) content encoders to match warm collaborative representations, while knowledge distillation (10.1145/3696410.3714852) uses content features as a medium to transfer teacher-model (warm-item) knowledge to student models (cold-item).
@@
To alleviate these issues, this paper proposes a novel item representation fusion and enhancement method based on semantic IDs, termed COINS. Specifically, inspired by OneSearch (chen2025onesearchpreliminaryexplorationunified): a recently deployed end-to-end generative framework that has achieved successful application in e-commerce search. We leverage its proposed RQ-OPQ encoding (endowed with rich hierarchical semantic and collaborative signals) to enhance cold-start item representations. First, for the coarser-grained RQ encoding, we design an adaptive transfer mechanism to enable bidirectional transfer between high-frequency collaborative signals embedded in item IDs and dense semantic information within RQ encodings. Second, for the residual vectors of items obtained via RQ hierarchical quantization (i.e., OPQ encodings), which encapsulate the discriminative information of different items, we integrate such discriminative information into item representations through contrastive learning. Ultimately, the resulting item representations not only incorporate abundant content and collaborative information but also accurately capture the uniqueness of individual items, as illustrate in Figure 1 (c). The main contributions are as follows:
@@
  • We conduct comprehensive offline experiments and rigorous online A/B test, whose results fully validate the superiority of our proposed COINS method.
@@ 2.1. Cold Start for Search
Typically, the core of the cold-start problem lies in exploring how to predict user interaction behaviors for new items. In this work, we focus on the click-through rate (CTR) prediction task and investigate how to improve the performance of cold-start tasks through the synergy between Semantic IDs generated by large search models and collaborative signals. In this section, we first formally define the cold-start problem, then elaborate on the proposed COINS framework in detail. Formally, given a user U (user profile $U_{p}$), the user’s current search query Q, and the user’s historical behavior sequence $H_{U}$, cross features $C$, each sample in the CTR prediction task can be represented as $I=(id,X)$, which includes a label $y\in\{0,1\}$ (indicating whether the sample is clicked by the user): id denotes the random hash ID of the item, and $X$ represents the multi-dimensional feature set of the item (including content features, metadata features, etc.). The goal of cold-start CTR estimation is to learn a discriminative model $f$ to output the click probability of the sample:
@@ 2.3. Semantic Ids Enhanced Item Representation
In this section, we elaborate in detail on COINS, a cold-start item representation enhancement framework based on semantic IDs. This framework comprises two core modules: the first is an adaptive transfer and alignment mechanism between RQ encoding and item IDs, which aims to learn coarse-grained representations with both semantic information and collaborative signals; the second is an item discriminative information learning module based on OPQ encoding, through which fine-grained discriminative information is injected into item representations, as illustrated in Figure 2.
Figure 2. The framework of COINS.
@@ 3.2.3. Baselines
Since our proposed COINS framework is based on semantic ids to enhance the representation of cold start items, we have selected some models that are also based on semantic ids or cold start as our baseline:
SPM_SID (DBLP:conf/recsys/SinghV0KSZHHWTC24): hashing sub fragments of semantic ids sequences to adaptively obtain the semantics of different items.
DAS (DBLP:journals/corr/abs-2508-10584): Through joint training of semantic collaborative models, use contrastive learning to synchronously optimize the quantization and alignment process of semantic ids and item id.
SaviorRec (yao2025saviorrec): A cold start model based on multimodal information and semantic ids collaborative training to enhance item representation.
@@ 3.3. Offline Performance
Table 1. The offline performance of COINS.
@@
COINS 0.8725 0.6394 0.8741 0.6405 0.8528 0.6301
To verify the effectiveness of COINS in the click-through rate (CTR) prediction task, this section compares the performance of COINS with existing state-of-the-art methods (SPM_SID, DAS, SaviorRec) through offline experiments, as shown in Table 1. In the full-sample (All) scenario, COINS achieves the best performance in both AUC (+0.38pp) and GAUC (+0.34pp). This proves that COINS’s overall ranking ability and resistance to user bias are significantly superior to baselines. In the warm item scenario, COINS still maintains a leading advantage, which indicates that even in scenarios with sufficient user interaction data, COINS can still optimize the representation quality of warm items and further improve prediction accuracy through the fusion of collaborative signals and semantic information in RQ-OPQ encodings. The cold-start item scenario is the core advantage area of COINS, The improvement margin is much higher than that in the full-sample and warm item scenarios. This result verifies the effectiveness of COINS’s design—by leveraging the discriminative information learning of OPQ encoding and the adaptive transfer of RQ encoding, COINS effectively solves the problem of scarce collaborative signals for cold-start items, making it the current State-of-the-Art (SOTA) method for cold-start CTR prediction tasks.
@@ 3.4. Ablation Study
Table 2. The ablation study resluts of COINS.
@@
COINS 0.8725 0.6394 0.8741 0.6405 0.8528 0.6301
To verify the function of each module in COINS, we conducted sufficient ablation experiments, and the experimental results are shown in the Table 2. Firstly, we directly replace the original random ID in base model by RQ-OPQ encodings (short for only sid). Although it achieves comparable overall performance to the base model, it exhibits inferior performance on warm items compared to the original model due to collaborative signal misalignment. Secondly, we attempted to preserve random IDs and treat SIDs as a type of ID feature (short for iid+SID). This variant has a slight improvement compared to the base model, as the introduction of SID brings additional information increment, but it is still far inferior to the COINS model. Finally, we ablated the two modules of COINS (short for iid+RQ, iid+OPQ). Both variants can achieve significant benefits compared to the base model. By integrating the advantages of these two variants, COINS simultaneously learns the discriminative information and convergent semantic information of items, ultimately achieving the optimal performance.
@@ 3.5. Online A/B Testing
Table 3. The Online A/B testing resluts of COINS.
@@
COINS +1.720% +2.230% +3.512% +9.639%
To verify the efficiency of COINS’s online platform, we compared it with an online ranking model in the e-commerce mall search scenario through rigorous A/B testing. We conducted a 14 day experiment under real online traffic, with 5 days being the AA period and the remaining 9 days being the AB period. The base group and exp group were randomly assigned 7% of the overall real online traffic. The experimental results are shown in the Table 3. Our experimental focus indicators are the buyers and order volume, and we also report the relevant indicators in the cold start scenario in the Table 3. Experiments have shown that COINS can increase the overall number of buyers by 1.72% and the order volume by 2.23%. Simultaneously, COINS significantly improves the cold start scenario (buyer number +3.512% and order volume +9.639%) in mall ecommerce search.
@@ 4. Conclusion
To address cold-start challenges in e-commerce, we propose COINS: it fuses item ID collaboration and semantics RQ via adaptive transfer, and enhances item differentiation with OPQ contrastive learning. Offline experiments and online A/B testing confirm its effectiveness (improving buyer +1.72%, order volume +2.23%). In future work, we will explore the fusion of multimodal semantics and collaborative signals to adapt to more complex cold-start scenarios.